OpenStudy (anonymous):

plz help !!! Write the standard equation for the circle. center (10, –6), r = 6

OpenStudy (akashdeepdeb):

This is the standard equation of a cricle! (x-x1)2 + (y-y1)2 = r2 So (x1,y1) = (10,-6) r = 6 Put in the values in the equation! Use the values and expand to get the circle equation! Getting me? :)

OpenStudy (anonymous):

ok would this be correct ? (x + 10)2 + (y – 6)2 = 6 ...?

OpenStudy (anonymous):

@AkashdeepDeb

OpenStudy (akashdeepdeb):

6^2 = 36 Yes it is RIGHT! Now expand by usign the basic algebraic identites! And you'd get the equation of that circle? Got it? :)

OpenStudy (anonymous):

ye but i have to leave it like that to b a standard equation

OpenStudy (akashdeepdeb):

Oh Okay!! :D You are right! :)

OpenStudy (anonymous):

ty!

OpenStudy (akashdeepdeb):

:)